# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset (GIST)

### v4

@@ Appendix C Additional Experiment Results
We present comprehensive evaluation results using BLEU (Papineni et al., 2002), ChrF (Popović, 2015), ChrF++ (Popović, 2017), TER (Snover et al., 2006), and COMET (Rei et al., 2020) for translation into the five target languages across both evaluation sets from Table 10 to Table 61. We also present one good example and one bad example for the prompting-powered refinement and the word alignment method across all languages, as shown from Figure 13 to Figure 17.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 22.38 23.75 23.02 23.58
Bleu score std 24.43 25.89 24.92 25.48
Bleu 1gram precision mean 22.48 23.85 23.13 23.67
Bleu 1gram precision std 24.47 25.92 24.96 25.51
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 70.18 70.31 70.33 70.22
Bleu ref len std 35.62 35.60 35.84 35.67
Bleu pred len mean 70.08 70.24 70.17 70.25
Bleu pred len std 35.40 35.41 35.62 35.57
Bleu bp mean 0.99 0.99 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 22.87 24.24 23.51 24.08
Chrf score std 24.56 26.02 25.04 25.61
Ter score mean 91.12 89.53 90.37 89.78
Ter score std 29.39 31.08 29.94 30.63
Chrfpp score mean 22.87 24.24 23.51 24.08
Chrfpp score std 24.56 26.02 25.04 25.61
Table 10: BLEU, ChrF, ChrF++, and TER results for Arabic on the 60-60 evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 23.13 24.28 23.10 25.48
Bleu score std 24.72 25.84 24.57 27.26
Bleu 1gram precision mean 23.19 24.35 23.18 25.53
Bleu 1gram precision std 24.72 25.85 24.59 27.27
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 69.56 69.79 69.84 69.72
Bleu ref len std 35.32 35.54 35.56 35.55
Bleu pred len mean 69.64 69.89 69.83 69.89
Bleu pred len std 35.11 35.31 35.30 35.37
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 23.67 24.83 23.64 26.03
Chrf score std 24.92 26.04 24.78 27.44
Ter score mean 90.27 88.94 90.29 87.50
Ter score std 29.63 31.00 29.45 32.62
Chrfpp score mean 23.67 24.83 23.64 26.03
Chrfpp score std 24.92 26.04 24.78 27.44
Table 11: BLEU, ChrF, ChrF++, and TER results for Arabic on the 60-60 evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 23.58 24.64 23.57 24.08
Bleu score std 26.63 27.85 26.40 27.16
Bleu 1gram precision mean 23.62 24.69 23.64 24.13
Bleu 1gram precision std 26.64 27.86 26.43 27.18
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 70.40 70.59 70.67 70.61
Bleu ref len std 35.88 35.95 36.14 36.11
Bleu pred len mean 70.68 70.85 70.88 70.87
Bleu pred len std 35.92 36.00 36.14 36.14
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.01 1.00 1.00 1.01
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 24.06 25.13 24.06 24.56
Chrf score std 26.76 27.98 26.55 27.28
Ter score mean 89.66 88.49 89.54 89.12
Ter score std 32.09 33.43 31.74 32.65
Chrfpp score mean 24.06 25.13 24.06 24.56
Chrfpp score std 26.76 27.98 26.55 27.28
Table 12: BLEU, ChrF, ChrF++, and TER results for Arabic on the 60-60 evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 19.98 20.52 19.77 20.56
Bleu score std 22.35 23.49 22.14 23.37
Bleu 1gram precision mean 20.03 20.57 19.82 20.61
Bleu 1gram precision std 22.37 23.50 22.15 23.40
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 69.72 69.93 69.91 69.91
Bleu ref len std 35.74 36.02 35.93 35.83
Bleu pred len mean 69.99 70.20 70.14 70.21
Bleu pred len std 35.69 35.97 35.85 35.79
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 20.50 21.03 20.28 21.08
Chrf score std 22.54 23.65 22.34 23.54
Ter score mean 93.97 93.40 94.20 93.36
Ter score std 26.96 28.24 26.69 28.17
Chrfpp score mean 20.50 21.03 20.28 21.08
Chrfpp score std 22.54 23.65 22.34 23.54
Table 13: BLEU, ChrF, ChrF++, and TER results for Arabic on the 60-60 evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 20.11 21.34 20.29 21.36
Bleu score std 23.09 24.68 23.43 24.39
Bleu 1gram precision mean 20.20 21.43 20.38 21.43
Bleu 1gram precision std 23.14 24.73 23.49 24.44
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 71.54 71.47 71.70 71.36
Bleu ref len std 36.10 36.29 36.28 36.33
Bleu pred len mean 71.79 71.75 71.93 71.68
Bleu pred len std 36.24 36.43 36.39 36.47
Bleu bp mean 0.99 0.99 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.01
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 20.62 21.85 20.80 21.87
Chrf score std 23.26 24.83 23.60 24.54
Ter score mean 93.61 92.27 93.37 92.26
Ter score std 27.75 29.58 28.12 29.27
Chrfpp score mean 20.62 21.85 20.80 21.87
Chrfpp score std 23.26 24.83 23.60 24.54
Table 14: BLEU, ChrF, ChrF++, and TER results for Arabic on the 60-60 evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 17.29 19.21 18.31 19.69
Bleu score std 22.88 24.37 24.26 25.91
Bleu 1gram precision mean 17.37 19.30 18.40 19.74
Bleu 1gram precision std 22.94 24.43 24.32 25.96
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 24.74 26.52 24.75 26.10
Bleu ref len std 12.69 13.49 12.72 13.40
Bleu pred len mean 24.59 26.35 24.60 26.09
Bleu pred len std 12.46 13.22 12.49 13.29
Bleu bp mean 0.99 0.99 0.99 1.00
Bleu bp std 0.02 0.02 0.02 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 17.45 19.36 18.47 19.84
Chrf score std 23.08 24.54 24.43 26.07
Ter score mean 83.26 81.33 82.23 80.86
Ter score std 23.01 24.53 24.40 26.07
Chrfpp score mean 17.45 19.36 18.47 19.84
Chrfpp score std 23.08 24.54 24.43 26.07
Table 15: BLEU, ChrF, ChrF++, and TER results for Chinese on the 60-60 evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 26.26 27.23 27.06 29.01
Bleu score std 24.84 26.18 25.59 27.21
Bleu 1gram precision mean 26.51 27.50 27.32 29.21
Bleu 1gram precision std 24.89 26.23 25.65 27.26
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 28.70 29.35 28.76 29.13
Bleu ref len std 14.62 14.70 14.65 14.61
Bleu pred len mean 28.24 28.86 28.30 28.85
Bleu pred len std 13.91 13.97 13.96 14.16
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.03 0.03 0.03 0.03
Bleu ratio mean 0.99 0.99 0.99 1.00
Bleu ratio std 0.04 0.04 0.04 0.04
Chrf score mean 26.43 27.41 27.23 29.17
Chrf score std 24.92 26.25 25.66 27.27
Ter score mean 74.32 73.34 73.51 71.57
Ter score std 25.21 26.56 25.97 27.60
Chrfpp score mean 26.43 27.41 27.23 29.17
Chrfpp score std 24.92 26.25 25.66 27.27
Table 16: BLEU, ChrF, ChrF++, and TER results for Chinese on the 60-60 evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 32.64 34.24 33.30 34.12
Bleu score std 29.93 31.10 30.52 30.34
Bleu 1gram precision mean 32.69 34.30 33.34 34.16
Bleu 1gram precision std 29.94 31.12 30.53 30.35
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 29.97 30.02 29.94 29.97
Bleu ref len std 14.74 14.79 14.77 14.75
Bleu pred len mean 30.02 30.06 30.00 30.02
Bleu pred len std 14.73 14.74 14.77 14.72
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 32.76 34.36 33.42 34.24
Chrf score std 29.94 31.11 30.53 30.35
Ter score mean 67.79 66.19 67.13 66.31
Ter score std 30.20 31.39 30.79 30.62
Chrfpp score mean 32.76 34.36 33.42 34.24
Chrfpp score std 29.94 31.11 30.53 30.35
Table 17: BLEU, ChrF, ChrF++, and TER results for Chinese on the 60-60 evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 26.08 26.55 26.47 26.53
Bleu score std 25.69 26.49 25.63 26.37
Bleu 1gram precision mean 26.17 26.64 26.56 26.62
Bleu 1gram precision std 25.73 26.53 25.67 26.40
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 29.33 29.36 29.31 29.37
Bleu ref len std 14.47 14.48 14.52 14.49
Bleu pred len mean 29.25 29.28 29.23 29.28
Bleu pred len std 14.24 14.24 14.29 14.22
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 26.20 26.66 26.59 26.65
Chrf score std 25.72 26.52 25.66 26.40
Ter score mean 74.40 73.94 74.02 73.95
Ter score std 25.99 26.77 25.91 26.64
Chrfpp score mean 26.20 26.66 26.59 26.65
Chrfpp score std 25.72 26.52 25.66 26.40
Table 18: BLEU, ChrF, ChrF++, and TER results for Chinese on the 60-60 evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 27.31 28.63 27.55 27.84
Bleu score std 27.56 28.76 27.73 28.04
Bleu 1gram precision mean 27.47 28.75 27.71 27.96
Bleu 1gram precision std 27.61 28.80 27.78 28.10
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 30.61 30.47 30.57 30.51
Bleu ref len std 14.86 14.93 14.87 14.98
Bleu pred len mean 30.54 30.46 30.50 30.51
Bleu pred len std 14.81 14.83 14.82 14.88
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.04 0.02 0.04 0.02
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.04 0.03 0.04 0.03
Chrf score mean 27.52 28.83 27.76 28.04
Chrf score std 27.69 28.88 27.86 28.16
Ter score mean 73.47 72.14 73.22 72.94
Ter score std 27.92 29.14 28.10 28.42
Chrfpp score mean 27.52 28.83 27.76 28.04
Chrfpp score std 27.69 28.88 27.86 28.16
Table 19: BLEU, ChrF, ChrF++, and TER results for Chinese on the 60-60 evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 34.93 37.78 35.14 38.16
Bleu score std 29.70 31.33 29.69 31.98
Bleu 1gram precision mean 35.10 37.91 35.31 38.28
Bleu 1gram precision std 29.76 31.37 29.75 32.02
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 100.59 101.31 101.12 101.26
Bleu ref len std 48.91 49.80 49.06 49.16
Bleu pred len mean 100.42 101.37 100.88 101.26
Bleu pred len std 48.75 49.74 48.84 49.02
Bleu bp mean 0.99 1.00 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 35.55 38.40 35.76 38.76
Chrf score std 29.85 31.44 29.83 32.08
Ter score mean 75.21 71.97 74.95 71.49
Ter score std 34.89 36.82 34.81 37.49
Chrfpp score mean 35.55 38.40 35.76 38.76
Chrfpp score std 29.85 31.44 29.83 32.08
Table 20: BLEU, ChrF, ChrF++, and TER results for French on the 60-60 evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 40.04 42.12 39.47 41.72
Bleu score std 31.35 32.61 31.61 32.25
Bleu 1gram precision mean 40.17 42.22 39.60 41.84
Bleu 1gram precision std 31.38 32.63 31.64 32.29
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 101.49 101.82 101.87 102.01
Bleu ref len std 50.11 50.22 50.16 50.25
Bleu pred len mean 101.56 101.99 101.88 102.13
Bleu pred len std 49.98 50.12 49.98 50.11
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 40.68 42.76 40.10 42.35
Chrf score std 31.39 32.64 31.65 32.26
Ter score mean 69.19 66.81 69.85 67.29
Ter score std 36.78 38.18 37.00 37.78
Chrfpp score mean 40.68 42.76 40.10 42.35
Chrfpp score std 31.39 32.64 31.65 32.26
Table 21: BLEU, ChrF, ChrF++, and TER results for French on the 60-60 evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 40.80 43.88 41.30 43.18
Bleu score std 32.29 33.65 32.31 33.31
Bleu 1gram precision mean 40.90 43.98 41.41 43.27
Bleu 1gram precision std 32.30 33.67 32.32 33.33
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 102.43 102.86 102.68 102.85
Bleu ref len std 49.81 50.12 49.94 50.07
Bleu pred len mean 102.54 102.98 102.72 102.97
Bleu pred len std 49.77 50.05 49.84 49.98
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 41.40 44.46 41.89 43.77
Chrf score std 32.27 33.61 32.28 33.28
Ter score mean 68.32 64.76 67.77 65.56
Ter score std 37.85 39.39 37.85 39.02
Chrfpp score mean 41.40 44.46 41.89 43.77
Chrfpp score std 32.27 33.61 32.28 33.28
Table 22: BLEU, ChrF, ChrF++, and TER results for French on the 60-60 evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 33.85 36.13 33.74 36.33
Bleu score std 28.54 30.46 28.57 30.77
Bleu 1gram precision mean 33.98 36.23 33.87 36.44
Bleu 1gram precision std 28.57 30.48 28.60 30.79
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 100.78 101.15 101.19 101.29
Bleu ref len std 49.88 49.82 49.92 49.92
Bleu pred len mean 100.70 101.22 101.07 101.32
Bleu pred len std 49.64 49.64 49.61 49.73
Bleu bp mean 0.99 1.00 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 34.48 36.74 34.37 36.94
Chrf score std 28.63 30.53 28.66 30.82
Ter score mean 76.35 73.71 76.53 73.46
Ter score std 33.54 35.72 33.57 36.11
Chrfpp score mean 34.48 36.74 34.37 36.94
Chrfpp score std 28.63 30.53 28.66 30.82
Table 23: BLEU, ChrF, ChrF++, and TER results for French on the 60-60 evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 33.05 35.50 33.24 35.02
Bleu score std 28.46 30.64 28.72 30.28
Bleu 1gram precision mean 33.18 35.62 33.38 35.15
Bleu 1gram precision std 28.50 30.67 28.75 30.32
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 102.37 102.52 102.66 102.71
Bleu ref len std 49.92 50.07 49.96 50.03
Bleu pred len mean 102.37 102.62 102.62 102.77
Bleu pred len std 49.76 49.92 49.77 49.85
Bleu bp mean 0.99 1.00 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 33.68 36.12 33.86 35.64
Chrf score std 28.56 30.71 28.81 30.36
Ter score mean 77.43 74.65 77.20 75.15
Ter score std 33.50 36.00 33.75 35.54
Chrfpp score mean 33.68 36.12 33.86 35.64
Chrfpp score std 28.56 30.71 28.81 30.36
Table 24: BLEU, ChrF, ChrF++, and TER results for French on the 60-60 evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 6.19 8.61 6.72 9.14
Bleu score std 12.90 16.15 13.59 17.19
Bleu 1gram precision mean 6.26 8.67 6.79 9.18
Bleu 1gram precision std 12.94 16.17 13.64 17.20
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 36.22 38.30 36.51 38.18
Bleu ref len std 20.17 20.98 20.20 20.76
Bleu pred len mean 35.45 37.76 35.74 37.85
Bleu pred len std 19.73 20.62 19.75 20.74
Bleu bp mean 0.98 0.98 0.98 0.99
Bleu bp std 0.03 0.03 0.03 0.03
Bleu ratio mean 0.98 0.99 0.98 0.99
Bleu ratio std 0.04 0.04 0.04 0.03
Chrf score mean 6.22 8.64 6.75 9.17
Chrf score std 12.93 16.17 13.62 17.20
Ter score mean 94.32 91.91 93.79 91.38
Ter score std 13.12 16.39 13.81 17.42
Chrfpp score mean 6.22 8.64 6.75 9.17
Chrfpp score std 12.93 16.17 13.62 17.20
Table 25: BLEU, ChrF, ChrF++, and TER results for Japanese on the 60-60 evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 14.56 15.30 14.61 16.94
Bleu score std 20.15 20.98 19.70 22.63
Bleu 1gram precision mean 14.71 15.43 14.76 17.02
Bleu 1gram precision std 20.30 21.10 19.86 22.68
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 40.61 40.93 40.74 41.08
Bleu ref len std 19.85 20.05 19.91 20.14
Bleu pred len mean 40.24 40.66 40.37 40.98
Bleu pred len std 19.70 19.97 19.76 20.10
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.02 0.02 0.02 0.02
Bleu ratio mean 0.99 0.99 0.99 1.00
Bleu ratio std 0.03 0.03 0.03 0.02
Chrf score mean 14.65 15.39 14.70 17.02
Chrf score std 20.22 21.05 19.77 22.67
Ter score mean 86.02 85.28 85.97 83.65
Ter score std 20.37 21.21 19.92 22.89
Chrfpp score mean 14.65 15.39 14.70 17.02
Chrfpp score std 20.22 21.05 19.77 22.67
Table 26: BLEU, ChrF, ChrF++, and TER results for Japanese on the 60-60 evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 21.46 22.11 21.66 22.40
Bleu score std 23.81 24.77 23.93 25.26
Bleu 1gram precision mean 21.47 22.11 21.66 22.40
Bleu 1gram precision std 23.81 24.77 23.93 25.26
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 43.57 43.54 43.59 43.57
Bleu ref len std 21.17 21.10 21.08 21.15
Bleu pred len mean 43.88 43.85 43.90 43.88
Bleu pred len std 21.21 21.14 21.11 21.18
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.00 0.00 0.00 0.00
Bleu ratio mean 1.01 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 21.59 22.24 21.79 22.53
Chrf score std 23.89 24.85 24.02 25.34
Ter score mean 79.33 78.68 79.13 78.39
Ter score std 24.23 25.20 24.36 25.69
Chrfpp score mean 21.59 22.24 21.79 22.53
Chrfpp score std 23.89 24.85 24.02 25.34
Table 27: BLEU, ChrF, ChrF++, and TER results for Japanese on the 60-60 evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 15.06 15.94 15.43 16.31
Bleu score std 18.77 19.65 18.65 20.25
[…]

### v5

@@ D.1 Additional Quantitative Results
We present comprehensive evaluation results using ChrF (Popović, 2015), ChrF++ (Popović, 2017), and TER (Snover et al., 2006) for translation into the five target languages across both evaluation sets in Table 15. By comparing the fourth and second values in each column of Table 2, we observe that the post-hoc prompting ablation enhances translation scores, aligning them with the performance of the prompting-powered method across all languages and models.
We also present one good example and one bad example for the prompting-powered refinement and the word alignment method across all languages, as shown from Figure 14 to Figure 18.
D.2 Qualitative Results
We also conduct a qualitative manual evaluation of the generation results produced by constrained beam search and logit adjustment methods, as explained in Section C.3. Both approaches exhibit extremely slow performance, running approximately 100 times slower than post-hoc methods. Furthermore, forcing specific word ids to appear in the output did not perform well, suffering from similar issues as the word alignment method in disrupting syntactical dependencies and morphological agreements. During manual inspection, we observe that constrained beam search behaves similarly to the logit adjustment method with a large scaling factor: the generated sentences often failed to maintain proper grammar, and terminological terms were frequently repeated multiple times within the same output. Conversely, with a small scaling factor in the logit adjustment method, the terms were often omitted entirely, demonstrating little to no effect on the output.
D.3 Statistical Tests for Experiment Results
To assess the statistical significance of our findings in Section 6.3, we conducted one-sided paired t-tests on the BLEU, COMET, ChrF, ChrF++, and TER scores. Based on our observations, we formulated the following three hypotheses:
  1. $H_{0}^{(1)}$:
    The prompting-powered refinement method outperforms direct translation across all languages.
  2. $H_{0}^{(2)}$:
    The word alignment method outperforms direct translation for Chinese and Japanese but underperforms for Arabic, French, and Russian.
  3. $H_{0}^{(3)}$:
    The prompting-powered refinement method outperforms the word alignment method across all languages.
As shown in Table 14, hypotheses 1 and 3 are fully supported by all metrics across all languages. Hypothesis 2 is partially supported, as there are cases where we cannot reject the null hypothesis that the word alignment method underperforms default translation for Arabic, French, and Russian. This outcome is insightful, as it suggests that the default translation approach remains effective for these languages.
Arabic Chinese French Japanese Russian
Metric Hypothesis T Stats P Val T Stats P Val T Stats P Val T Stats P Val T Stats P Val
COMET 1 14.18 0.00 15.22 0.00 12.68 0.00 11.13 0.00 8.53 0.00
2 6.91 0.00 7.60 0.00 1.84 0.03 3.21 0.00 20.76 0.00
3 17.93 0.00 11.67 0.00 13.58 0.00 10.62 0.00 24.89 0.00
BLEU 1 6.37 0.00 6.72 0.00 9.80 0.00 6.26 0.00 7.63 0.00
2 -0.66 0.75 3.89 0.00 -0.22 0.59 2.58 0.00 -2.85 1.00
3 4.80 0.00 3.87 0.00 8.77 0.00 4.57 0.00 5.50 0.00
ChrF 1 6.37 0.00 6.72 0.00 9.79 0.00 6.26 0.00 7.64 0.00
2 -0.66 0.75 3.89 0.00 -0.18 0.57 2.59 0.00 -2.85 1.00
3 4.79 0.00 3.86 0.00 8.76 0.00 4.55 0.00 5.49 0.00
ChrF++ 1 6.37 0.00 6.72 0.00 9.79 0.00 6.26 0.00 7.64 0.00
2 -0.66 0.75 3.89 0.00 -0.18 0.57 2.59 0.00 -2.85 1.00
3 4.79 0.00 3.86 0.00 8.76 0.00 4.55 0.00 5.49 0.00
TER 1 6.17 0.00 6.72 0.00 9.71 0.00 6.26 0.00 7.41 0.00
2 -0.85 0.80 3.88 0.00 -0.18 0.57 2.59 0.00 -2.76 1.00
3 4.40 0.00 3.87 0.00 8.73 0.00 4.56 0.00 5.27 0.00
Table 14: Hypothesis test statistics and p-values for all metrics across the five tested models.
Model Metric Arabic Chinese French Japanese Russian
Evaluation Set: 60-60
ChrF 20.62 + 1.24 + 0.18 + 1.26 27.52 + 1.32 + 0.24 + 0.52 33.68 + 2.44 + 0.18 + 1.96 14.76 + 0.61 + 0.32 + 1.08 16.99 + 1.59 - 0.06 + 0.92
[…]